# Understanding Self-Regulated Learning Behavior Among High and Low Dropout Risk Students During CS1: Combining Trace Logs, Dropout Prediction and Self-Reports

### v1

@@ 4.3.3. Active task-focused learners
The active task-focused learner profile type ($N_{\text{student}}=5$) represents students who focused primarily on completing the weekly tasks. Unlike the two aforementioned profiles, these students consistently employed effective, low dropout risk strategies throughout the course, with only one student having initial adjustment struggles during the first week. The active focus on course tasks can be seen in low dropout predictions of weekly strategies and the high median percentage of completed weekly tasks. The active task-centered approach is reflected in the self-reports, exemplified by the following quote:
@@
However, high task activity does not guarantee a high grade. One student mentions language barrier as an external obstacle for their performance:
@@
[…]
@@ 4.3.4. Persevering until the end
A notable outlier to all students is Student 1, who completed the course with the best grade despite only minimally interacting with weekly tasks. Even when relying only on high dropout risk strategies, the student achieved the best grade by obtaining near best points from the final exam. While the final grade can be seen an indicator of good performance, the student’s self-evaluation of own learning methods is much more critical, with them setting a low bar for own performance:
@@
Visually, the student’s strategy was generally focused on a quick and risky attempt to complete mandatory tasks, and in course week 9 there is are no observes student sessions at all. The risky approach to studying is confirmed in the self-report:
@@
When answering the question about scheduling learning, the student notes their reactive nature to deadlines:
@@
Further, the student studied alone, highlighting avoidance of social events as the main reason:
@@
Nevertheless, the main reason to passing the course could be attributed to the student’s ability to find and information necessary:
@@ 4.3.5. Course dropouts
The final profile type ($N_{\text{student}}=7$) includes all participants who dropped out during the course. While no students in the profile filled a self-report questionnaire, the Brunner-Munzel tests confirm that both grades ($\hat{p}^{*}(39)=27.221,p<.001,\hat{p}^{\prime\prime}=0.975$) and dropout predictions ($\hat{p}^{*}(20.362)=14.408,p<.001,\hat{p}^{\prime\prime}=0.085$) of the students in the profile in are not stochastically equal to the other students. Evaluating the effect size, the students in this profile tend to have lower grades and higher predicted dropout than the other students.
@@
  • four students dropped out during the first course weeks with only high dropout risk strategies used;
  • two managed to hang on the course until the course difficulty considerably increased in weeks 3 and 5;
  • one student started with low dropout risk strategies and managed to continue through initial complex topics, however they quickly fell to using high dropout risk strategies and dropped out right after a particularly difficult topic was introduced.
In other words, even if a student in this profile did manage to use low dropout risk strategies, ultimately every student reverted to high-risk strategies 1–2 weeks prior to dropout. In this profile, all students eventually dropped out by the midway of the course when programming tasks became more complex.
@@ 5. Discussion
This study aimed to examine the behavioral patterns of CS1 students throughout the course by analyzing trace logs and student dropout predictions, interpreting the results through the lens of SRL. The findings demonstrate that combining event-based logs, dropout predictions, and basic self-reports reveals clear differences between high and low dropout risk CS1 student behaviors, providing detailed insights into their successes and struggles throughout the course. Although detecting and understanding learning patterns associated with different dropout risks can help address course attrition, research into automatic pattern detection remains scarce in CSEd (Omer et al., 2023). This study therefore contributes to the research area by demonstrating a method for semi-automated learning pattern analysis that complements and extends qualitative inquiries into learning patterns of programming students, such as those by Liao et al. (2019). Furthermore, by relating trace log data to self-reports to obtain a more comprehensive narrative of student behavior during CS1, this study contributes to the discussions regarding practical applications of event-based measures (e.g., Moreno-Marcos et al., 2020; Fan et al., 2023; Han, 2023; Choi et al., 2023).
@@ 5.1. Addressing the research questions
Regarding RQ1, we identified three clusters among strategies with low dropout probability and interpreted them as the following weekly learning strategy types: task-oriented, prioritizing main task performance, seeking understanding, and resource-focused, prioritizing course materials over tasks. These strategies align with Liao et al. (2019), who reported that high-performing students tend to use more varied course resources. Furthermore, information seeking and experimentation with simpler tasks before solving more complex problems are common SRL strategies in computer science education, as exemplified by Falkner et al. (2014) and Silva et al. (2024). Although prioritizing task completion is not generally reported as an SRL strategy itself, it can be considered a form of goal-setting and planning, as students allocate more time and effort to tasks (e.g., Garcia et al., 2018). Additionally, from clustering student behaviors over the entire course, we identified three clusters that we interpreted as learning profile types associated with course success: adaptive understanding seekers, diligent course content followers, and active task-focused learners. These profile types relate to each other in that most of the students of these types applied low dropout risk strategies throughout the course, and often actively varying strategies depending on the weekly topic. This aligns with Liao et al. (2019), who noted that high-performing students tend to vary their study approaches depending on the task.
@@
Regarding RQ2, we identified nine distinct clusters among strategies with high dropout probability from which we interpreted high dropout risk strategy types. Unlike the low dropout risk strategy types, these strategy types show greater variation in behavior and dropout risk. For instance, one of the interpreted strategy types, slow start, finding study pace, does not by itself indicate dropout, as study participants who employed this strategy improved their learning style as the course progressed. This strategy type aligns with the common observation that students may initially struggle to self-regulate in a new context but learn it over time (e.g., Loksa and Ko, 2016; Winne and Marzouk, 2019). Most of the interpreted strategy types, such as attempting with examples; low engagement; late reliance on model answers; and struggling to understand, needing help, represent in our view basic temporary unsuccessful strategies that even successful students may experience and recover from, as observed in the identified student profile clusters (cf. Figure 5). Such notion has been noted as well by Liao et al. (2019), who noted that lower-performing students can overcome problems just as high-performing students may temporarily struggle. However, the final interpreted strategy types — falling behind, realizing struggle too late; superficial review, stuck on a single resource; browsing materials with sporadic task attempts; and risky focus on mandatory tasks only — seem to describe behaviors of students on the verge of dropping out. In the case of the course dropouts present in this study, these strategy types were among the last ones those students attempted before dropping out. In our view, these strategy types can be related to issues with time management and difficulty assessment, both identified by Falkner et al. (2014) as the most frequent unsuccessful SRL strategies in CS. On the other hand, these strategy types can be simply seen as acts of giving up prior to dropout. In other words, those students who persevere through problems recovered from high-risk strategies, while dropouts remained locked in high-risk behavior until they were overwhelmed by the workload — a common narrative for dropping CS1 (Petersen et al., 2016; Kinnunen and Malmi, 2006).
@@
Regarding RQ3, we found that SRL as we observed and interpreted via trace logs and students’ self-reported evaluations of their own SRL are generally consistent and complementary. Although we initially mapped learning strategies and student profiles using only trace logs and dropout prediction, self-reports enhanced the interpretation of the types. Specifically, both trace log-based data and self-reports were equally valuable but provided different perspectives. While Liao et al. (2019) and our self-reports offer a snapshot of students’ cognition at the end of the course, trace logs allow following the full learning path of a student and capture their overall learning profile as transitions between learning strategies (cf. Figure 5). Although the somewhat open-ended self-report tool we used may not be appropriate for predicting dropout on its own, it aids describing student behavior. In this study, the self-reports were instrumental in being able to differentiate the three main learning profile types: adaptive understanding seekers, diligent course content followers, and active task-focused learners. Furthermore, the profile of a student who passed the course with full marks despite being identified as high-risk (see Section 4.3.4) would not have a reasonable explanation without the student’s self-report. Thus, our results confirm an argument made in an earlier study (Zhidkikh et al., 2023): using predictive learning analytics and trace logs alongside self-reports allows for a more comprehensive understanding of how and why a student uses the available learning resources. Aligning with numerous prior studies (Tempelaar et al., 2020; Zhidkikh et al., 2024; Fan et al., 2023; Lim et al., 2021), we posit that the inclusion of self-reports can improve the interpretation of purely trace log-based analyses by contextualizing them and bringing in cognitive aspects that cannot be easily captured as atomic events.
@@ 5.2. Practical implications
5.3. Limitations and threats to validity
We recognize several threats to validity in our study design. In this subsection, we have argued for our chosen research setting, its limitations, and how we have strived to mitigate potential threats to validity.
Relatively small number of participants: The number of participants in this study is relatively small considering the total number of students in the course, and considering that the data were analyzed with statistical methods. However, despite the number of participants in the statistical analyses (N = 47), the number of observations collected from 11 course weeks is higher, i.e., 5,445 student sessions grouped into 458 weekly learning strategies. Furthermore, while we used the Brunner-Munzel test mainly because our data failed to meet the assumption for independence of observations due to our study design, the Brunner-Munzel test has been shown more robust with smaller sample sizes than other adjacent non-parametric tests. Finally, SRL trace log analysis has been applied successfully on a relatively low number of participants before (Zhidkikh et al., 2023).
@@ 5.3. Limitations and threats to validity
Arbitrary number of clusters: Arguably, the data, as visualized in Figure 5, could be clustered with a higher or lower number of clusters, the former increasing the precision on which the results are reported, and the latter abstracting student behavior. We chose the number of clusters based on the observations of the first author, who is deeply familiar with the design of the course in this study. Based on these observations, a different numbers of clusters were initially considered, yet 12 clusters captured detailed enough behavior while providing a sufficient level of abstraction for the results to be relatively easy to interpret.
@@
Arbitrary study session length: We chose an arbitrary length of 25 minutes for a study session in the trace log analysis, which potentially skews the results in unforeseeable ways. We chose this number based on our previous observations on how much time it is expected for a student to complete one exercise. Some other studies (e.g., Jovanović et al., 2017; Matcha et al., 2019) have used a 30 minute cutoff, which is also arbitrary. Other heuristics could have potentially been more accurate, but there are no standardized rules for determining study sessions from trace logs (Kovanovic et al., 2015). Reliably determining when a study session ends would have required a more controlled experiment (e.g., a laboratory setting, or capturing participant screens). We hypothesized that such a requirement would have yielded considerably less participants.
@@
Self-report answer interpretation: In addition to the statistical tests, we analyzed the self-reports of 25 participants using content analysis. While the results yielded by content analysis are subject to researcher interpretation and subjective bias, we strived to mitigate this by having two authors analyze the data independently, discussing possible discrepancies, and forming a consensus, as detailed in Section 3.2.2. Despite this approach, another group of researchers might have formed at least slightly different set of themes.
Single-cohort study: The participants were recruited from a single cohort in a single university course. This approach poses potential challenges to the generalizability of the results concerning, e.g., topics besides CS1 or computing in general, cross-cultural differences, different teaching approaches, or some intra-cohort intricacies not considered in this study. As with many other research settings, these potential differences may be considered in replication studies.
@@ 6. Conclusion
In this study, we investigated the behavioral patterns of students in a CS1 course and their interactions with the course’s learning resources. By applying trace log analysis and student dropout prediction methods developed in two prior studies (Zhidkikh et al., 2023, 2024) to a cohort of CS1 students, we identified and interpreted 12 weekly learning strategy types based on students’ interactions with course resources and four core learning profile types that describe how these interactions change over the course duration. Among these learning strategy types, three were interpreted from low dropout risk behaviors and nine from high risk behaviors, indicating that dropout risk can be observed in various forms and may require different types of interventions. We compared the interpreted strategy types with other studies on self-regulated learning, finding close connections to related works such as Liao et al. (2019). Additionally, our observations of student behavior prior to dropout was consistent with prior studies that map how students drop CS1 (e.g., Kinnunen and Malmi, 2006; Petersen et al., 2016). Altogether, previous studies provide initial support and validation to our interpretations.
@@
Nevertheless, this study has limitations that warrant further studies. Most importantly, different CS1 contexts and with larger cohorts is needed. Given that CS1 course we obtained participants from had 321 enrolled students, there may be more unique learning paths that our identified strategy types and learning profile types do not fully encompass. Different CS1 course contexts may reveal additional nuances in student behavior, particularly concerning dropout. Additionally, we see potential for a reproduction study (cf. Ihantola et al., 2015, , Figure 2) by examining paths to dropout using longitudinal analysis tools. For example, applying Latent Transition Analysis (Lanza et al., 2003) to study sessions could uncover more intricate learning paths leading to dropout.

### v2

@@ 5.5.3. Active task-focused learners
The active task-focused learner profile type ($N_{\text{student}}=5$) represents students who focused primarily on completing the weekly tasks. Unlike the two aforementioned profiles, these students consistently employed effective, low dropout risk strategies throughout the course, with only one student showing initial adjustment struggles during the first week. That student did not specify the reason behind the slower start and mentioned only a generally slower learning pace:
[Student 27]: […] even though I have slipped through my schedule a bit, I was able to get the course assignment and weekly exercises done well.
Overall, the active focus on course tasks is reflected in the low dropout predictions of the weekly strategies and the high median percentage of completed weekly tasks. The task-centered approach is also reflected in the self-reports (3 students), exemplified by the following quote:
@@
However, high task activity does not guarantee a high grade. One student mentions a language barrier as an external obstacle to their performance:
@@
By visual inspection, the active task-focused learners’ self-reports appear more positive than those of the other students. However, the overall self-reported opinions of students in this profile type appear stochastically equal to those of the other self-report answerers (permuted Brunner-Munzel, $p=.713$). While the median grade appears higher, there is no statistically significant support for the difference (permuted Brunner-Munzel, $p=.350$). The dropout predictions of the strategies used by students in this profile type are not stochastically equal to those of the other study participants ($\hat{p}^{*}(73.978)=-4.379,p<.001,\hat{p}^{\prime\prime}=0.662$). Based on the effect size, students in this profile tended to use more low dropout risk strategies than the other study participants, agreeing with the visualization of the learning strategy types used.
@@ 5.5.4. Persevering until the end
A notable outlier among all students is Student 1, who completed the course with the highest grade despite only minimally interacting with the weekly tasks. Even though they relied only on high dropout risk strategies, the student achieved the top grade by obtaining near-best points on the final exam. While the final grade can be seen as an indicator of good performance, the student’s self-evaluation of their own learning methods is much more critical, setting a low bar for their own performance:
@@
Visually, the student’s strategy was focused on a quick, risky attempt to complete the mandatory tasks, and in course week 9 there are no observed student sessions at all. This risky approach to studying is confirmed in the self-report:
@@
When answering the question about scheduling learning, the student notes a reactive approach to deadlines:
@@
Further, the student studied alone, citing avoidance of social events as the main reason:
@@
Nevertheless, the main reason for passing the course could be attributed to the student’s ability to find the necessary information:
@@ 5.5.5. Course dropouts
The final profile type ($N_{\text{student}}=7$) includes all participants who dropped out during the course. While no students in this profile completed the self-report questionnaire, Brunner-Munzel tests confirm that both the grades ($\hat{p}^{*}(39)=27.221,p<.001,\hat{p}^{\prime\prime}=0.975$) and the dropout predictions ($\hat{p}^{*}(20.362)=14.408,p<.001,\hat{p}^{\prime\prime}=0.085$) of the students in this profile are not stochastically equal to those of the other students. Based on the effect sizes, the students in this profile tend to have lower grades and higher predicted dropout than the other students.
@@
  • four students dropped out during the first course weeks, using only high dropout risk strategies;
  • two held on to the course until the difficulty increased considerably in weeks 3 and 5;
  • one student started with low dropout risk strategies and managed to continue through the initial complex topics but quickly shifted to high dropout risk strategies and dropped out right after a particularly difficult topic was introduced.
In other words, even when a student in this profile managed to use low dropout risk strategies, all students reverted to high-risk strategies 1–2 weeks prior to dropout. All students in this profile eventually dropped out by the midpoint of the course, when the programming tasks became more complex.
@@ 6. Discussion
This study aimed to examine the behavioral patterns of CS1 students throughout the course by analyzing trace logs and student dropout predictions, interpreting the results through the lens of SRL. The findings demonstrate that combining event-based logs, dropout predictions, and basic self-reports reveals clear differences between high and low dropout risk CS1 student behaviors, providing detailed insights into their successes and struggles throughout the course. Although detecting and understanding learning patterns associated with different dropout risks can help address course attrition, research into automatic pattern detection remains scarce in CSEd (Omer et al., 2023). This study therefore contributes to the research area by demonstrating a method for semi-automated learning pattern analysis that complements and extends qualitative inquiries into learning patterns of programming students, such as those by Liao et al. (2019). Furthermore, by relating trace log data to self-reports to obtain a more comprehensive narrative of student behavior during CS1, this study contributes to the discussions regarding practical applications of event-based measures (e.g., Moreno-Marcos et al., 2020; Fan et al., 2023b; Han, 2023; Choi et al., 2023).
@@ 6.1. Addressing the research questions
Regarding the learning behaviors associated with low dropout risk (RQ1), we identified three clusters among strategies with low dropout probability and interpreted them as the following weekly learning strategy types: task-oriented, prioritizing main task performance; seeking understanding; and resource-focused, prioritizing course materials over tasks. These strategies align with Liao et al. (2019), who reported that high-performing students tend to use more varied course resources. Furthermore, information seeking and experimentation with simpler tasks before solving more complex problems are common SRL strategies in computer science education, as exemplified by Falkner et al. (2014) and Silva et al. (2024). Although prioritizing task completion is not generally reported as an SRL strategy itself, it can be considered a form of goal-setting and planning, as students allocate more time and effort to tasks (e.g., Garcia et al., 2018). Additionally, from clustering student behaviors over the entire course, we identified three clusters that we interpreted as learning profile types associated with course success: adaptive understanding seekers, diligent course content followers, and active task-focused learners. These profile types relate to each other in that most of the students of these types applied low dropout risk strategies throughout the course, and often actively varying strategies depending on the weekly topic. This aligns with Liao et al. (2019), who noted that high-performing students tend to vary their study approaches depending on the task.
@@
Regarding the learning behaviors associated with high dropout risk (RQ2), we identified nine distinct clusters among strategies with high dropout probability from which we interpreted high dropout risk strategy types. Unlike the low dropout risk strategy types, these strategy types show greater variation in behavior and dropout risk. For instance, one of the interpreted strategy types, slow start, finding study pace, does not by itself indicate dropout, as study participants who employed this strategy improved their learning style as the course progressed. This strategy type aligns with the common observation that students may initially struggle to self-regulate in a new context but learn it over time (e.g., Loksa and Ko, 2016; Winne and Marzouk, 2019). Most of the interpreted strategy types, such as attempting with examples; low engagement; late reliance on model answers; and struggling to understand, needing help, represent in our view basic temporary unsuccessful strategies that even successful students may experience and recover from, as observed in the identified student profile clusters (cf. Figure 6). Such notion has been noted as well by Liao et al. (2019), who noted that lower-performing students can overcome problems just as high-performing students may temporarily struggle. However, the final interpreted strategy types — falling behind, realizing struggle too late; superficial review, stuck on a single resource; browsing materials with sporadic task attempts; and risky focus on mandatory tasks only — seem to describe behaviors of students on the verge of dropping out. In the case of the course dropouts present in this study, these strategy types were among the last ones those students attempted before dropping out. In our view, these strategy types can be related to issues with time management and difficulty assessment, both identified by Falkner et al. (2014) as the most frequent unsuccessful SRL strategies in CS. On the other hand, these strategy types can be simply seen as acts of giving up prior to dropout. In other words, those students who persevere through problems recovered from high-risk strategies, while dropouts remained locked in high-risk behavior until they were overwhelmed by the workload — a common narrative for dropping CS1 (Petersen et al., 2016; Kinnunen and Malmi, 2006).
@@
Regarding the interplay between the self-reports and trace log data (RQ3), we found that SRL as we observed and interpreted via trace logs and students’ self-reported evaluations of their own SRL are generally consistent and complementary. Although we initially mapped learning strategies and student profiles using only trace logs and dropout prediction, self-reports enhanced the interpretation of the types. Specifically, both trace log-based data and self-reports were equally valuable but provided different perspectives. While Liao et al. (2019) and our self-reports offer a snapshot of students’ cognition at the end of the course, trace logs allow following the full learning path of a student and capture their overall learning profile as transitions between learning strategies (cf. Figure 6). Although the somewhat open-ended self-report tool we used may not be appropriate for predicting dropout on its own, it aids describing student behavior. In this study, the self-reports were instrumental in being able to differentiate the three main learning profile types: adaptive understanding seekers, diligent course content followers, and active task-focused learners. Furthermore, the profile of a student who passed the course with full marks despite being identified as high-risk (see Section 5.5.4) would not have a reasonable explanation without the student’s self-report. Thus, our results confirm an argument made in an earlier study (Zhidkikh et al., 2023a): using predictive learning analytics and trace logs alongside self-reports allows for a more comprehensive understanding of how and why a student uses the available learning resources. Aligning with numerous prior studies (Tempelaar et al., 2020b; Zhidkikh et al., 2024b; Fan et al., 2023b; Lim et al., 2021), we posit that the inclusion of self-reports can improve the interpretation of purely trace log-based analyses by contextualizing them and bringing in cognitive aspects that cannot be easily captured as atomic events.
@@ 6.2. Practical implications
6.3. Trustworthiness and limitations
We recognize several considerations regarding trustworthiness and limitations in our study design. In this subsection, we have argued for our chosen research setting, its limitations, and how we have strived to mitigate potential threats to validity.
Number of participants: The number of student participants in this study is a potential limitation considering the total number of students in the course. The sample size poses potential threats to validity with regard to sampling bias, as participation was voluntary and we cannot ensure a truly random sample of CS1 students. However, despite the limited number of participants (N = 47), the number of observations relevant to the study is larger: 5,445 student sessions grouped into 458 weekly learning strategies, collected over 11 weeks of the course. The number of participants is apparent primarily in the characterization of students’ study profiles, where statistical analysis was used as an additional aid rather than as the primary method of characterization. Because our data did not meet the independence assumption due to the study design, we used the Brunner-Munzel test for statistical comparison. This test has been shown to be robust with smaller sample sizes than other comparable non-parametric tests. SRL trace log analysis has also been applied successfully with a relatively small number of participants in earlier work (Zhidkikh et al., 2023a).
@@ 6.3. Trustworthiness and limitations
Selection of appropriate clustering: The clustering performed in Section 5.2 and Section 5.4 could be done with a higher or lower number of clusters, yielding clustering at different levels of granularity. To support our choice of clustering, we computed multiple cluster validation indices as a guide. However, each index tends to be sensitive to different aspects of the clustering (Liu et al., 2010; Schubert, 2023). Cluster analysis is ultimately an exploratory method that can yield multiple valid and interesting solutions, and relying solely on statistical indices may not be enough to produce a meaningful result (Schubert, 2023; Salvador and Chan, 2004). Our choice of clustering was therefore informed by the indices, by our domain knowledge of the studied CS1 course, and by the pragmatic need to capture a sufficient level of abstraction for the results to be relatively easy to interpret. To ensure transparency, we reported our observations and our process for choosing the cluster numbers in the sections referenced above.
@@
Selection of study session length: In Section 4.1, we set a 25-minute length for a study session in the trace log analysis, which affected the number and the granularity of the study sessions. We chose this value based on the course context and our estimate of how much time a student would typically need to complete one exercise. Other heuristics could have been more accurate, but there are no standardized rules for determining study sessions from trace logs (Kovanovic et al., 2015). Some other studies (e.g., Jovanović et al., 2017; Matcha et al., 2019a) have used a 30-minute cut-off without an explicit rationale or grounding in the course context. Reliably determining when a study session ends would have required a more controlled experiment (e.g., a laboratory setting or capturing participant screens). We hypothesized that such a requirement would have yielded considerably less participants.
@@
Self-report answer interpretation: In addition to the statistical tests, we analyzed the self-reports of 29 participants using content analysis. While the results yielded by content analysis are subject to researcher interpretation and subjective bias, we strived to mitigate this by having two authors analyze the data independently, discussing possible discrepancies, and forming a consensus, as detailed in Section 4. Despite this approach, another group of researchers might have formed at least slightly different set of themes.
Nature of self-reports: A potential concern is that self-reports, by nature, are prone to recall bias. In this study, self-reports were administered before the course exam, while students’ memory of their study habits was still fresh. Further, self-reports were used as only one of several data sources, with trace logs providing objective information on student performance. We did not observe any discrepancies between the two data sources that would suggest recall bias, although we acknowledge that some self-reports may omit information that could affect the interpretation of the student profiles. Another potential source of bias is social desirability, i.e., the tendency of participants to respond in a favorable way. In this study, this is mitigated by the voluntary nature of participation, which provided no direct benefit to the students.
Single-cohort study: The participants were recruited from a single cohort in a single university course. This approach poses potential challenges to the generalizability of the results concerning, e.g., topics besides CS1 or computing in general, cross-cultural differences, different teaching approaches, or some intra-cohort intricacies not considered in this study. For example, the studied CS1 course allowed students to compensate for low task performance with good exam results, yielding the Persevering until the end profile that might not emerge under other course designs. However, self-regulation is generally dependent on the course context (Ben-Eliyahu and Bernacki, 2015), and the syllabus, teaching methods, and assessment practices can all affect the specific approaches students take to learning. From a methodological standpoint, we posit that grand generalizations of student behavior are difficult to make without assuming a closed system and absolute control over all external factors, which are considered problematic (Bhaskar, 1978, p. 67). Even with a different cohort and sample size, the behavior patterns identified in such a study would not negate those identified in the present cohort. We therefore emphasize the exploratory nature of our results and the value of the analysis pipeline to the course context in which the method is applied. Finally, as with many other research settings, these potential differences may be examined in replication studies.
@@ 7. Conclusion
In this study, we investigated the behavioral patterns of students in a CS1 course and their interactions with the course’s learning resources. By applying trace log analysis and student dropout prediction methods developed in two prior studies (Zhidkikh et al., 2023a, 2024b) to a cohort of CS1 students, we identified and interpreted 12 weekly learning strategy types based on students’ interactions with course resources and four core learning profile types that describe how these interactions change over the course duration. Among these learning strategy types, three were interpreted from low dropout risk behaviors and nine from high risk behaviors, indicating that dropout risk can be observed in various forms and may require different types of interventions. We compared the interpreted strategy types with other studies on self-regulated learning, finding close connections to related works such as Liao et al. (2019). Additionally, our observations of student behavior prior to dropout was consistent with prior studies that map how students drop CS1 (e.g., Kinnunen and Malmi, 2006; Petersen et al., 2016). Altogether, previous studies provide initial support and validation to our interpretations.
@@
Nevertheless, this study has limitations that warrant further research. Most importantly, studies in different CS1 contexts and with larger cohorts are needed. Given that the CS1 course from which we recruited participants had 321 enrolled students, there may be additional learning paths that our identified strategy types and learning profile types do not fully capture. Different CS1 contexts may also reveal further nuances in student behavior, particularly concerning dropout. Additionally, we see potential for a reproduction study (cf. Ihantola et al., 2015, , Figure 2) by examining paths to dropout using longitudinal analysis tools. For example, applying Latent Transition Analysis (Lanza et al., 2003) to study sessions could uncover more intricate learning paths leading to dropout.